### v1

# Defining Boundaries: A Spectrum of Task Feasibility for Large Language Models

Abstract
Large language models (LLMs) have shown remarkable performance in various tasks but often fail to handle queries that exceed their knowledge and capabilities, leading to incorrect or fabricated responses. This paper addresses the need for LLMs to recognize and refuse infeasible tasks due to the required skills surpassing their capabilities. We first systematically conceptualize infeasible tasks for LLMs, providing formal definitions and categorizations that cover a spectrum of related hallucinations. We develop and benchmark a new dataset comprising diverse infeasible and feasible tasks to test multiple LLMs’ abilities on task feasibility. Furthermore, we explore the potential of training enhancements to increase LLMs’ refusal capabilities with fine-tuning. Experiments validate the effectiveness of our methods, offering promising directions for refining the operational boundaries of LLMs in real applications.
Defining Boundaries: A Spectrum of Task Feasibility for
Large Language Models
Wenbo Zhang   Zihang Xu   Hengrui Cai University of California Irvine {wenbz13,zxu18,hengrc1}@uci.edu
**footnotetext: Equal Contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding Author
1 Introduction
Large language models (LLMs) have made significant breakthroughs in addressing diverse tasks (Brown et al., 2020; Wei et al., 2022; Chowdhery et al., 2023). One primary concern with LLMs lies in their dishonesty or hallucinations in handling queries beyond their knowledge and capabilities. Ideally, when LLMs lack the relevant knowledge, they should either decline to respond or indicate uncertainty. Yet, often, they generate incorrect or fabricated information, leading to undesirable erroneous outputs. A few recent studies have been proposed on these issues. Liu et al. (2024) introduced the UnknownBench benchmark to evaluate how well various LLMs can express uncertainty in scenarios where they lack adequate parametric knowledge. Similarly, studies by Amayuelas et al. (2023) and Yin et al. (2023) explore how LLMs distinguish between queries within and beyond their knowledge scopes. Additional works (Yang et al., 2023; Zhang et al., 2023a; Cheng et al., 2024) aim to align LLMs to acknowledge their own limitations, prompting them to state "I don’t know" when faced with unfamiliar questions. However, all these studies mainly assess the models’ hesitance to refuse responses that surpass their knowledge with a focus on the question-answering tasks. A broader examination of what LLMs can and cannot handle, i.e., their general capabilities, is thus in demand.
Figure 1: Illustration example: given an infeasible instruction (requiring physical interaction), a desirable LLM is expected to refuse the query but the undesirable LLM will be reluctant to refuse and generate incorrect or irrelevant responses (hallucinations).
Table 1: Four main categories for infeasible tasks with concise descriptions, examples, and previous works that define unfeasible tasks.
Category Description Example Previous Work
Physical Interaction Physical interaction and execution of actions in the real world "Change my car tire on the side of the road" N/A
Virtual Interaction Interaction with digital environments or external virtual tool "Book a flight for me next month"; "Who is the winner of World Cup 2026" Yang et al. (2023); Sun et al. (2024); Liu et al. (2024); Cheng et al. (2024)
Non-text Input or Output Process or create non-text data "Translate spoken language in a video into another language" Sun et al. (2024)
Self-awareness Recognize and understand oneself as a distinct entity "Explain an instance where you surprised others with your actions" N/A
Real-world applications usually involve tasks beyond simple factual question answering (Sun et al., 2024), such as text summarization, ticket booking, online information retrieval, etc. These tasks demand a variety of skills, and we deem a task infeasible for LLMs if it requires skills that exceed the capabilities of language models. For instance, as shown in Fig. 1, suppose we request an LLM with the query "Please dust the bookshelf in the living room"; a desirable model is expected to either decline to respond or express low confidence, as such a physical task falls outside the operational scope of a language model. This leads to a fundamental question of LLMs’ hallucination: are LLMs capable of expressing uncertainty or choosing not to respond when they lack the necessary skills?
To comprehensively examine and answer this question, in this paper, we focus on text-to-text language models that operate independently of external tools since this is the fundamental backbone of current advanced multimodal LLMs (Wu et al., 2023; Liu et al., 2023; Li et al., 2023) and AI agents (Schick et al., 2024; Shen et al., 2024). We first formally define infeasible tasks for LLMs and categorize them into four types: 1. Physical Interaction. 2. Virtual Interaction. 3. Non-text Input or Output. 4. Self-awareness. Our study is broad in scope and encompasses previous research that discusses tasks deemed infeasible as shown in Table 1. For example, when LLMs lack up-to-date knowledge to answer questions (see e.g., Yang et al., 2023; Sun et al., 2024; Liu et al., 2024; Cheng et al., 2024), it belongs to our second category - Virtual Interaction - since online information querying is required. Utilizing the proposed definitions, we can further generate benchmark data (see details in Fig. 2) that exemplify these infeasible tasks. Additionally, we assemble a set of feasible tasks to serve as control groups in our study. The primary objective of this study is to determine whether current state-of-the-art LLMs can accurately differentiate between feasible and infeasible tasks when provided with specific definitions.
With the definition of task feasibility, we are also interested in whether additional training can enhance the refusal capabilities of LLMs for infeasible tasks without relying on explicit prompting. Traditional supervised fine-tuning approaches (see e.g., Ouyang et al., 2022; Wang et al., 2022b) typically force models to generate completed outputs. Consequently, these models attempt to provide answers even when confronted with queries beyond their abilities. Recent research (Zhang et al., 2023a; Cheng et al., 2024) indicates that training models only on correct responses may inadvertently condition them to speculate instead of acknowledge their limitations. This observation motivates us to develop a new training approach using a dataset augmented with refusal responses to infeasible tasks. By doing so, we aim to fine-tune models with appropriate abilities to decline infeasible queries. We explore multiple strategies to construct such a training dataset to enhance its effectiveness.
Our contributions to this field are threefold:
$\bullet$ We are the first study to systematically conceptualize tasks that are infeasible for LLMs, providing a formal definition and categorization of these tasks. Our work covers a spectrum of hallucinations related to task feasibility.
$\bullet$ We establish a new dataset for task feasibility, comprising a diverse range of commonly posed infeasible and feasible tasks, and benchmark multiple LLMs under the developed dataset, providing valuable insights for future research in this area.
$\bullet$ We propose three strategies to enhance the refusal awareness of LLMs when faced with infeasible tasks, by constructing a refusal-augmented instruction tuning dataset. Extensive experimental results demonstrate the effectiveness of these strategies.
Figure 2: Dataset constructing pipeline for the Infeasible Benchmark. It includes three stages: 1. Infeasible Task Definition. 2. Infeasible/feasible Task Generation. 3. Post Process for cleaning the generated data.
2 Proposed: Infeasible Benchmark
[…]
2.1 Define Infeasible Tasks
Infeasible tasks for LLMs refer to requests or queries that fall outside the operational scope or capabilities of these models. Commonly characterized as out-of-distribution (OOD), these tasks often demand actions or outputs that LLMs are not designed to handle. For instance, LLMs cannot perform physical actions like taking photographs or executing real-world tasks such as cooking. Additionally, these models might struggle with highly specialized knowledge not covered during their training or scenarios requiring real-time data updates, such as stock market analysis. Thus, recognizing and managing infeasible or out-of-distribution tasks is crucial for effectively utilizing LLMs and setting realistic expectations for their performance.
To develop a comprehensive definition and description of infeasible tasks, we review existing datasets that identify OOD tasks and conduct thorough human inspections of each example. Additionally, to deepen our understanding of what constitutes feasible versus infeasible tasks for LLMs, we deep dive into a series of structured inquiries with an LLM. Specifically, we posed questions such as, "What are feasible and infeasible tasks for large language models?" followed by a set of clarifying questions to refine the responses. To ensure the reliability of our findings, this questioning process was repeated multiple times, primarily utilizing the GPT-4 model (OpenAI, 2023).
Finally, based on the previous datasets (Sun et al., 2024; Zheng et al., 2023; Zhang et al., 2023b) and collected information from GPT-4, we categorize infeasible tasks into four main categories:
1. Physical Interaction: These are tasks that require direct physical interaction with the real world. They involve the manipulation of or interaction with physical objects or environments, such as moving items, operating machinery, or physically engaging with various materials.
2. Virtual Interaction: This category includes tasks that necessitate interaction within digital or virtual environments. These tasks may involve navigating web interfaces, utilizing virtual tools like search engines to gather new information, or executing commands within software applications.
3. Non-text Input or Output: These tasks involve dealing with data in formats other than text, such as images, audio, video, and sensory data.
4. Self-awareness: These tasks require an understanding of self-awareness, where the entity recognizes and comprehends its existence as a separate, sentient individual capable of introspection and self-reflection.
Our definitions constitute an exhaustive categorization of infeasible tasks that align closely with findings from prior research (see e.g., Yang et al., 2023; Sun et al., 2024; Liu et al., 2024; Cheng et al., 2024), effectively acting as a superset. For each category, we include illustrative examples and pertinent references, as detailed in Table 1.
2.2 Automatic Data Generation
Our objective is to develop a dataset that encompasses a wide range of queries with limited manual intervention. Inspired by the self-instruct methodology (Wang et al., 2022a), initially, we establish a small seed set of manually crafted tasks, which serve to direct the subsequent generation process. Subsequently, we prompt the model to formulate instructions for novel tasks, utilizing the example tasks from the seed set to facilitate the creation of tasks with broader coverage. Additionally, we enhance the prompts with formal task definitions, as this has been observed to yield more accurate and satisfactory generative outcomes. We also generate feasible tasks as a control group. We use a similar prompt by replacing the example tasks. The prompting templates for generating data are shown in Appendix E.
2.3 Quality Check
During the filtering stage, we employ Sentence-BERT (Reimers and Gurevych, 2019) to automatically evaluate each question source. We establish a similarity threshold of $0.97$, an empirically determined value aimed at effectively removing questions with excessive similarity. This is supplemented by a manual quality review to further eliminate any duplicate or ambiguous entries. Our analyses indicate that the text length of generated feasible data typically exceeds that of infeasible data. To facilitate a fair comparison between feasible and infeasible datasets, we divide the generated data into three distinct length categories. Within these categories, we conduct a one-to-one matching to standardize the length distribution across both datasets.
The final benchmark dataset is composed of feasible and infeasible parts. For the infeasible part, each of the four categories comprises $25\%$ of the total. Summary statistics of our benchmark are in Appendix A.
3 Distinguish Feasible and Unfeasible Tasks with Uncertainty Scores
Utilizing the proposed Infeasible Benchmark, we aim to evaluate various strategies for expressing uncertainty to determine their effectiveness in distinguishing between feasible and unfeasible tasks. Considering the application of these strategies in both open-source and closed-source models, we focus on verbalized confidence elicitation. This approach involves prompting LLMs to explicitly articulate the reliability of their responses in natural language. This is particularly vital for closed-source models, which restrict interactions to text input-output and do not provide access to token logits (Lin et al., 2022; Xiong et al., 2023). In this study, we employ a regression-style method of elicitation, where LLMs provide confidence scores on a scale from 0 to 100, reflecting their perceived accuracy of the response.
3.1 Evaluation Setup
Methods. Here we utilize four types of verbalized confidence methods. All methods require the LLM to output a confidence score that the given instruction is feasible without answering the instruction but in different ways of querying LLMs.
$\bullet$ Pre-response: directly ask for the confidence score without answering the instruction.
$\bullet$ Mid-response: first identify and classify the category of the given instruction and then ask for the confidence score.
$\bullet$ Post-response: first answer the given instruction and then ask for the confidence score.
$\bullet$ Mix-response: combination of mid and post-response.
Pre-response is the simplest way of getting the confidence score. Mid, Post, and Mix-response let the LLM have more thinking steps before outputting the final score. The prompting templates for each method are shown in Appendix E.
Models. we conduct a collection of experiments with GPT-3.5 (February 2024 version), GPT-4 (April 2024 version), PaLM2 (April 2024 version) (Anil et al., 2023), and the chat version of LlaMA2-70b (Touvron et al., 2023).
Table 2: Measuring distinguishability and calibration for various models and methods. Bold number represents the best one for each individual model. We also do a cross-model comparison and represent the best-$4$ methods for each metric. It can be seen that GPT-4 archives the best performance for all metrics, showing its superior ability to recognize feasible tasks.
Model Method Metric
AUROC ()($\uparrow$)( ↑ ) KSS ()($\uparrow$)( ↑ ) Brier Score ()($\downarrow$)( ↓ )
LLaMA2-70b-chat Pre $\mathbf{0.927}$ $\mathbf{0.723}$ $\mathbf{0.107}$
Mid $0.896$ $0.688$ $0.131$
Post $0.914$ $0.718$ $0.119$
Mix $0.841$ $0.570$ $0.191$
PaLM2 pre $\mathbf{0.913}$ $\mathbf{0.725}$ $0.111$
Mid $0.898$ $0.696$ $0.123$
Post $0.910$ $0.716$ $0.115$
Mix $0.896$ $0.667$ $\mathbf{0.132}$
GPT-3.5-turbo Pre $0.858$ $0.575$ $0.173$
Mid $0.865$ $\mathbf{0.633}$ $0.167$
Post $0.855$ $0.540$ $0.188$
Mix $\mathbf{0.886}$ $0.622$ $\mathbf{0.150}$
GPT-4 Pre $0.965$ $\mathbf{0.892}$ $\mathbf{0.056}$
Mid $0.955$ $0.884$ $0.061$
Post $\mathbf{0.967}$ $0.878$ $0.061$
Mix $\mathbf{0.967}$ $0.880$ $\mathbf{0.056}$
Table 3: Measuring distinguishability and calibration for various models and methods for long-form instructions. Bold number represents the best one for each individual model.
Model Method Metric
AUROC ()($\uparrow$)( ↑ ) KSS ()($\uparrow$)( ↑ ) Brier Score ()($\downarrow$)( ↓ )
LLAMA2-70b-chat Pre 0.672 0.280 0.272
Mid 0.550 0.159 0.375
Post 0.542 0.153 0.375
Mix 0.549 0.229 0.351
PaLM2 Pre 0.562 0.123 0.934
Mid 0.778 0.504 0.198
Post 0.514 0.027 0.499
Mix 0.514 0.027 0.496
GPT-3.5-turbo Pre 0.770 0.396 0.269
Mid 0.693 0.291 0.328
Post 0.605 0.370 0.369
Mix 0.657 0.242 0.277
GPT-4 Pre 0.865 0.753 0.141
Mid 0.849 0.636 0.177
Post 0.859 0.643 0.180
Mix 0.810 0.554 0.204
Metrics. We evaluate distinguishability using two metrics: the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Kolmogorov–Smirnov Statistic (KSS). The AUROC measures the probability that a model ranks a randomly selected positive instance higher than a randomly selected negative instance. An AUROC value of 1.0 signifies perfect classification accuracy, whereas a value of 0.5 indicates no better performance than random guessing. The KSS assesses the maximum distance between the cumulative distribution functions of two sets of samples, with higher values indicating greater separation between distributions. In addition, we assess model calibration, which examines the correspondence between a model’s expressed confidence and its actual accuracy. "We selected the Brier Score as our metric because it favors probability predictions that are both well-calibrated and precise, aligning with our purpose.
Table 4: GPT-4 can still distinguish feasible and unfeasible data when the dataset generation was switched to GPT-3.5.
Method AUROC KSS Brier Score
Pre 0.941 0.805 0.091
3.2 Results and Analyses
Table 2 presents the outcomes of various methods used to derive confidence scores from different LLMs. We provide a summary of several critical insights from these experiments. $\mathbf{1}$. Excluding GPT-3.5, the pre-response method generally outperforms other models. This suggests that adding additional thinking steps does not typically enhance performance for LLMs. This finding is particularly notable, as in prior studies—like differentiating between ambiguous and unambiguous questions (Hou et al., 2023) or identifying known versus unknown queries (Liu et al., 2024)—increased reasoning steps (e.g., mid, post, and mix-response) proved advantageous. A plausible explanation is that for more advanced LLMs, the process of identifying feasible versus unfeasible instructions becomes more straightforward. $\mathbf{2}$. Across all models and techniques, GPT-4 consistently delivers the most precise and well-calibrated confidence estimates through direct verbalization compared to other models, which is also shown in Fig. 3. Additionally, GPT-4 exhibits minimal variability in results across different methods; for instance, the AUROCs for pre and post are 0.965 and 0.967, respectively.
To further validate our results on more complex and real scenarios, we also create an additional benchmark dataset focused on long instructions, where each instruction comprises multiple tasks. The results in Table 3 indicate that long-form instructions are more challenging for current LLMs to accurately determine their feasibility compared to short-form benchmarks. For instance, GPT-4 using the pre-method achieved an AUROC of only $0.865$, significantly lower than the $0.965$ achieved in the previous short-form benchmark. Also, the overall calibration of probability becomes less well-aligned, which might make the model outputs less trustworthy. Those results highlight the increased difficulty of processing long-form instructions.
Our benchmark dataset was initially generated using GPT-4. To assess the potential data leakage, we conducted an ablation study where the dataset generation was switched to GPT-3.5. Subsequently, we evaluated GPT-4 using this newly generated dataset ($n=400$). The results are presented in Table 4 and show that GPT-4 can robustly and consistently distinguish feasible and unfeasible data.
Figure 3: The Histogram of verbalized confidence from the pre-response method for 4 models. It can be seen that GPT-4 has the sharpest confidence in distinguishing feasible and infeasible data.
4 Can We Teach LLM to Refuse Unfeasible Tasks without Hints?
With the benchmark, we observed that state-of-the-art LLMs can differentiate between feasible and infeasible tasks when provided with carefully designed query prompts. However, in practical applications, users typically interact with LLMs using straightforward queries without complex instructions. This raises a fundamental question: can we train LLMs to autonomously reject infeasible tasks during routine interactions without extensive prompting?
Our findings indicate that when presented with questions that exceed their capabilities, LLMs tend to attempt an answer. This occurs because training models solely on feasible tasks inadvertently condition them to provide responses, rather than recognizing and communicating their limitations. If a model is not specifically trained to express "I can’t do this" as a valid response, it lacks the capability to do so when faced with infeasible tasks. To address this issue, we emphasize the importance of equipping a model to intelligently respond based on its inherent capabilities. Hence, this motivates us to refine our model to accurately express confidence levels and decline to execute infeasible instructions.
4.1 Methods
Given an instruction tuning dataset, we first reconstruct a refusal-added dataset where we explicitly incorporate refusal words into the data. We also find that some of the data belongs to infeasible tasks and might introduce hallucinations for refusal awareness. Here we have two strategies to achieve this.
4.1.1 Selection-based
We employ a two-stage training framework in our methodology. The initial phase focuses on identifying and recognizing data instances within the instruction-tuning dataset that are beyond the capability of the original model. Upon identifying these uncertain instances, we modify the dataset by substituting the original responses with refusal expressions for infeasible queries, while maintaining the original responses for feasible queries.
To enhance the diversity of refusal expressions, we crafted multiple variations of refusal text. These expressions are detailed in Appendix D. For the identified infeasible data, we employ random sampling to select appropriate refusal expressions. This approach ensures a varied and comprehensive response strategy for handling queries that exceed the model’s capabilities.
4.1.2 Augment-based
Rather than selecting uncertain data points, we incorporate new infeasible instruction data directly into the original dataset. For these newly added infeasible data points, we randomly select refusal expressions from a predefined set.
Table 5: Refusal rates and win rate of LLMs evaluated on our test dataset. The win rate is calculated in terms of LLaMA2-7b-chat.
Model Method OOD (Infeasible) Alpaca (Feasible)
Refusal Rate()($\uparrow$)( ↑ ) Win rate ()($\uparrow$)( ↑ ) Refusal Rate ()($\downarrow$)( ↓ )
OpenLLaMA-3b-v2 Original $0.105$ $0.357$ $\mathbf{0.059}$
Random $0.165$ $0.336$ $0.076$
Select $\mathbf{0.66}$ $0.335$ $0.086$
Augment $0.255$ $\mathbf{0.370}$ $0.069$
LLaMA2-7b Original $0.130$ $\mathbf{0.551}$ $0.070$
Random $0.140$ $0.296$ $0.184$
Select $\mathbf{0.735}$ $0.443$ $0.081$
Augment $0.175$ $0.432$ $\mathbf{0.065}$
LLaMA2-7b-chat $0.210$
GPT-3.5 $0.580$
GPT-4o $0.585$
4.1.3 Random-based
To underscore the significance of this selection process, we introduce an additional baseline, termed random-based, where queries are randomly chosen for response replacement. To ensure a fair comparison, we maintain the proportion of data undergoing response replacement consistent across different approaches.
Once the dataset has been augmented and structured, we proceed with standard supervised fine-tuning (SFT) on the newly constructed dataset.
4.2 Experimental Setting
Models. We use the Open-LLaMA-3B (Geng and Liu, 2023) and LLaMA-2-7b as the pre-trained model. We choose LLaMA-2-7b-chat as the reference model during the evaluation.
Metrics. We assess the models from two dimensions: helpfulness and refusal awareness. To evaluate helpfulness, we leverage recent advancements in automated evaluation, using a high-performing large language model, specifically GPT-4o, as a proxy for human labeling. In this evaluation, the model ranks pairs of responses, one generated by the trained model and the other by a reference model. We use the average win rate as the metric for this assessment. To mitigate position bias, responses are presented in both sequential orders, and the average rank is calculated. The prompting template for evaluation is shown in Appendix E.
For evaluating refusal awareness, we implement lexical keyword matching to calculate the refusal rate. This method involves identifying specific keywords that signify abstention, apology, or denial, enabling us to measure the model’s capacity to appropriately refuse or defer a response when necessary.
Data. Alpaca dataset (Taori et al., 2023) is a widely used instruction dataset and we use its cleaned version as our main training dataset. We split the original dataset into training and testing. To evaluate helpfulness, we utilize the test part of Alpaca. To evaluate the ability of refusal, we utilize an OOD dataset from Sun et al. (2024). More summary statistics for the datasets we used can be found in Appendix B.
To more comprehensively measure the general ability of fine-tuned models, we tested the model on another dataset with a larger size called Alpagasus, which is mentioned in Chen et al. (2023). This dataset contains $700$+ data, carefully curated from multiple resources, and is regarded as "feasible" to LLMs. Since the models we fine-tuned were trained using the Alpaca dataset, we consider this scenario as an evaluation of their ability to handle out-of-distribution data.
4.3 Experimental Results
We show our experiment results in Table 5 and summarise the main findings below.
LLMs without explicit refusal teaching exhibit limited refusal abilities: To see whether advanced LLMs can autonomously reject infeasible tasks without extensive prompting. We evaluate multiple advanced LLMs and find in general they exhibit limited refusal abilities. Even the best LLM (GPT-4o) among many benchmarks rejects only $58.1\%$ of the infeasible instructions, suggesting that refusal awareness is still lacking and additional explicit refusal teaching is necessary.
Selection matters to teach refusal: Among three methods of teaching refusal (Random, Select, and Augment), we find Select is the best way of increasing refusal awareness. It can help OpenLLaMA-3b-v2 and LLaMA2-7b achieve $66\%$ and $73.5\%$ respectively, which are far better than strong LLMs like GPT-4o, and GPT-3.5. We also can regard random can be seen as an ablation study of the selection step and we observe inferior results, showing the importance of the selection step.
When utilizing selection, we find there are approximately $7.5\%$ training data belonging to infeasible tasks. So if we replace their responses with refusal expressions, we can correctly incorporate the refusal instruction and reduce hallucination. On the contrary, for the augment method, the refusal rate is much lower, indicating that the augmentation with more infeasible data doesn’t eliminate the hallucination of the original dataset.
Trade-off between the helpfulness and refusal-awareness: We find this trade-off is similar to previous LLM studies (Bai et al., 2022; Touvron et al., 2023) when enhancing LLM’s instruction-following capabilities while ensuring they remain helpful and honest. We observe that there is a drop in general helpfulness. For example, in $3$b scale experiments, the win rate of select and random methods dropped nearly $2\%$ compared with original tuning (without refusal teaching). This is even worse with $7$b where all methods have over $10\%$ drop. This indicates that the proposed tunning methods still can’t provide a good balance between helpfulness and refusal-awareness.
Hard to improve general helpfulness: The results of testing the fine-tuned models on the Alpagasus dataset are shown in Table 9. The results show a general drop in win rate for all tunning methods compared with the original and suggest that these methods are not very resilient to distribution shifts and may not significantly improve general helpfulness. Therefore, future work should focus on developing more effective instruction-tuning methods to better manage distribution shifts.
5 Related Work
In this section, we review the progress on uncertainty quantification and hallucinations of large language models (LLMs).
5.1 Uncertainty Quantification in LLMs
Uncertainty quantification remains a core problem in deep learning. Guo et al. (2017) were among the first to point out that the predictive confidence of deep neural networks is often not well-calibrated. Recent studies have sought to address this by estimating and calibrating uncertainty specifically for language models (Xiao et al., 2022; Kuhn et al., 2023; Lin et al., 2023). A novel approach within this domain is verbalized confidence, which involves prompting LLMs to articulate their confidence levels in textual form Lin et al. (2022); Xiong et al. (2023). Tian et al. (2023) demonstrated that the method of verbalized confidence is effectively calibrated. Building on this straightforward approach, recent studies have further investigated its utility across various applications. These include tasks such as error detection (Xiao et al., 2022; Duan et al., 2023), ambiguity detection (Hou et al., 2023), and the identification of unanswerable queries (Liu et al., 2024). Our work can be seen as a generalization of utilizing the verbalized method in feasibility detection.
5.2 Hallucinations in LLMs
Despite the impressive performance characterized by high fluency and coherence, LLMs are still prone to generating unfaithful and nonfactual content, commonly referred to as hallucinations (Maynez et al., 2020). Several factors contribute to this phenomenon, including aspects of the training data, the algorithms used for training, and the inference processes (Ye et al., 2023; Zhang et al., 2023c; Rawte et al., 2023). Often, the training datasets themselves may include misinformation or become outdated, which can exacerbate the misalignment between the model’s outputs and factual accuracy (Penedo et al., 2023; Reddy et al., 2023; Li et al., 2024). Furthermore, LLMs have a tendency to overestimate their capabilities, leading them to produce incorrect responses with undue confidence and to struggle with recognizing when questions are unknown or unanswerable (Yin et al., 2023; Amayuelas et al., 2023; Cheng et al., 2024; Liu et al., 2024).
Recent research efforts have focused on addressing the issue of hallucinations in LLMs. For the detection of hallucinations, Azaria and Mitchell (2023) have developed a classifier that operates based on the internal states of LLMs. To measure the factuality of generated content, Lee et al. (2022) introduced a benchmark that utilizes both factual and nonfactual prompts. Furthermore, Varshney et al. (2023) employed an uncertainty-based approach to both detect and mitigate hallucinations during content generation. Zhang et al. (2023b) implemented a method that mimics human attention to factuality, guided by uncertainty scores. More recently, Sun et al. (2024) proposed out-of-distribution tasks but didn’t provide a formal definition and systematic summarization. There are also some recent recent works focused on investigating LLMs’ ability to abstain from answering to avoid hallucination (Slobodkin et al., 2023; Feng et al., 2024; Wen et al., 2024; Miyai et al., ). Our research contributes to this field by evaluating and training deliberate refusal of infeasible instructions, further aiding in the quantification and reduction of hallucinations in the era of LLMs.
6 Conclusion
We introduce the Infeasible Benchmark to analyze the behavior of LLMs when faced with instructions that exceed their capabilities. Our findings indicate that advanced LLMs are capable of distinguishing between feasible and infeasible tasks when provided with detailed guiding prompts. Yet, this capability diminishes in practical scenarios where users have minimal guidance regarding infeasible tasks. Additionally, we have developed fine-tuning methods aimed at enhancing the models’ refusal awareness. Our results show that the selection-based method demonstrates commendable performance in declining infeasible tasks.
Limitations
Despite the promising results of the proposed Infeasible Benchmark and fine-tuned models, we observe a trade-off between the helpfulness of responses and refusal awareness, suggesting that current approaches are not yet optimal. This identifies a clear avenue for future research. Our current definitions of feasibility are categorized at a coarse level into four groups. Future studies can introduce finer categorizations, which may enable more precise control over the behaviors of LLMs. Given that our research is limited to text-to-text language models, an intriguing direction for future work would be to extend the scope of infeasible task definitions to more advanced models, such as multimodal models or specific AI agents. This expansion could potentially aid in managing and controlling hallucinations more effectively. Additionally, another compelling exploration is to enhance refusal awareness while maintaining the level of helpfulness of these models.
Ethics Statement
This study focuses on providing formal definitions and categorizations of infeasible tasks of LLMs and a benchmark to access their identification. Our benchmark dataset is collected by querying GPT-4. Recognizing the ethical implications of using AI-generated data, we have implemented stringent measures to ensure the accuracy and reliability of the synthetic data while minimizing potential biases. We also assessed the ethical implications of deploying such a dataset, considering both its potential to innovate in the field and the necessity of mitigating any negative impacts on societal norms and individual privacy. This commitment underscores our dedication to responsible AI development and its application in linguistics.
Appendix A Benchmark Dataset Summary
Table 6: Summary statistics of Benchmark Dataset.
Feasible Infeasible
n 1850 1898
Length 10.04 9.47
Number for each category 430,531,464,473 1898
Table 7: Summary statistics of long instructions.
Feasible Infeasible
n 145 144
Length 44 60
Appendix B Instruction Tuning Dataset Summary
The dataset we used is under CC-BY-NC-4.0 license.
Table 8: Summary statistics of Instruction Dataset.
Alpaca OOD
# of train split 12784
# of test split 185 200
Table 9: Win rate and Refusal Rate of different models evaluated on additional test dataset Alpagasus.
Model Method Win rate ()($\uparrow$)( ↑ ) Refusal Rate ()($\downarrow$)( ↓ )
OpenLLAMA-3b-v2 Original 0.189 0.073
Random 0.176 0.081
Select 0.143 0.134
Augment 0.164 0.084
LLaMA2-7b Original 0.289 0.085
Random 0.149 0.133
Select 0.210 0.129
Augment 0.213 0.074
Appendix C Computing Resources
All of the computing experiments are conducted on a server with 4*A6000.
Appendix D Refusal Expression
It looks like I’m unable to complete the task you’re asking for.
Unfortunately, I’m not equipped to handle this specific request.
I’m sorry, but I can’t assist with that request.
I apologize, but what you’re asking is beyond my capabilities at the moment.
I can’t perform that action.
I’m sorry, but I cannot deal with this type of task.
I’m afraid I can’t take on this type of task.
I’m not equipped to deal with this.
Unfortunately, I can’t help with this particular task.
This seems a bit outside my scope.
Appendix E Prompt Used in Experiments
Figure 4: Prompt for generating infeasible data category 1: physical interaction.
Figure 5: Prompt for generating infeasible data category 2: virtual interaction.
Figure 6: Prompt for generating infeasible data category 3: non-text input or output.
Figure 7: Prompt for generating infeasible data category 4: self-awareness.
Figure 8: Prompt for generating feasible data.
Figure 9: Prompt for generating long feasible data.
Figure 10: Prompt for generating long infeasible data.
Figure 11: Prompt for pre-response method
Figure 12: Prompt for mid-response method
Figure 13: Prompt for post-response method
Figure 14: Prompt for mix-response method
Figure 15: Prompt for evaluating wint-rate.